OpenStudy (anonymous):

what is the next sequence in these 1,1,2,3,5,8 6,18,19,57,58 6,2,-4,-8-14

OpenStudy (anonymous):

In the first one you did 1+1=2, 2+3=5, 5+3=8, so now you must do 8+5=? I'm not sure about the second one or the third one.